Suppose we have a market demand Q = 18 – P and a cost C(Q) 9) = 3Q?.Suppose that firm 1 in the market described in question 1 has first mover advantage. (Market demand is Q 18 – P and both firm
Answer #1

First Mover Model (Stackelberg’s Model)

We have the following information

Market the demand function is

P = 18 – (Q1 + Q2)

Q = Q1 + Q2

and the duopolists' costs are

Total Cost of Firm 1: C1 = 0.5Q21

Total Cost of Firm 2: C2 = 0.5Q22

The reaction functions are found by taking the partial derivatives of the duopolists' profit functions and equating them to zero:

Π1 = PQ1 – C1

Π1 = [18 – (Q1 + Q2)]Q1 – 0.5Q21

Π1 = 18Q1 – Q21 – Q1Q2 – 0.5Q21

Π1 = 18Q1 – 1.5Q21 – Q1Q2

Π2 = PQ2 – C2

Π2 = [18 – (Q1 + Q2)]Q2 – 0.5Q22

Π2 = 18Q2 – Q22 – Q1Q2 – 0.5Q22

Π2 = 18Q2 – 1.5Q22 – Q1Q2

The partial derivatives are

∂Π1/∂Q1 = 18 – 3Q1 –Q2

∂Π2/∂Q2 = 18 – 3Q2 –Q1

The reaction functions are

Q1 = 6 – (1/3)Q2 ----- Firm 1 Reaction Curve

Q2 = 6 – (1/3)Q1 ------ Firm 2 Reaction Curve

Stackelberg's solution with Firm 1 being the sophisticated leader or first mover

Firm 1 will substitute the reaction function of Firm 2 in its own profit equation, which it will then maximise as if it were a monopolist:

Π1 = PQ1 – C1 = 18Q1 – 1.5Q21 – Q1Q2

Substitute, Q2 = Q2 = 6 – (1/3)Q1

Π1 = 18Q1 – 1.5Q21 – Q1(6 – (1/3)Q1)

Maximise:                                                       Π1 = 12Q1 – 1.17Q21

First-order condition: ∂Π1/∂Q1 = 12 – 2.34Q1 = 0

This yields output:                 Q1 = 5.1

Π1 = 12Q1 – 1.17Q21

Π1 = 12(5.1) – 1.17(5.1)2

Π1 = 30.77

Firm 2 would be the follower. It would assume that Firm 1 would produce 5.1 units; thus Firm 2 substitutes this amount in its reaction function

Q2 = 6 – (1/3)Q1

Q2 = 6 – (1/3)(5.1)

Q2 = 4.3

Π2 = 18Q2 – 1.5Q22 – Q1Q2

Π2 = (18 × 4.3) – 1.5(4.3)2 – (5.1 × 4.3)

Π2 = 77.4 – 27.74 – 21.93

Π2 = 27.73

Thus, the total output in the market is

Q = Q1 + Q2 = 5.1 + 4.3 = 9.4

And the market price

P = 18 – Q

P = 18 – 9.4

P = 8.6

Cournot Model

P = 18 – Q

P = 18 – (Q1 + Q2)

Cost function of Firm 1: C1 = 0.5Q21

Cost function of Firm 2: C2 = 0.5Q21

The profits of the duopolists are

Π1 = PQ1 – C1 = [18 – (Q1 + Q2)]Q1 – 0.5Q21

Π1 = 18Q1 – Q21 – Q1Q2 – 0.5Q21

Π1 = 18Q1 – 1.5Q21 – Q1Q2

Π2 = PQ2 – C2 = [18 – (Q1 + Q2)]Q2 – 0.5Q22

Π2 = 18Q2 – Q22 – Q1Q2 – 0.5Q22

Π2 = 18Q2 – 1.5Q22 – Q1Q2

For profit maximization under the Cournot assumption we have

∂Π1/∂Q1 = 0 = 18 – 3Q1 – Q2

∂Π2/∂Q2 = 0 = 18 – 3Q2 – Q1

The reaction functions are

Q1 = 6 – (1/3)Q2

Q2 = 6 – (1/3)Q1

Replacing Q2 into the Q1 reaction function we get

Q1 = 6 – 1/3[6 – (1/3)Q1]

Q1 = 4.5

And

Q2 = 6 – (1/3)Q1

Q2 = 6 – (1/3)(4.5)

Q2 = 4.5

Thus, the total output in the market is

Q = Q1 + Q2 = 4.5 + 4.5 = 9

And the market price

P = 18 – Q

P = 18 – 9

P = 9

Π1 = PQ1 – C1

Π1 = (9 × 4.5) – 0.5(4.5)2

Π1 = 40.5 – 10.125

Π1 = 30.36

Π2 = PQ2 – C2

Π2 = (9 × 4.5) – 0.5(4.5)2

Π2 = 40.5 – 10.125

Π2 = 30.36

In the First Mover Model the price is lower and output is higher as compared to the Cournot model. So, we can say that while consumers will prefer First Mover Model.

However, for Firm 2 profit is higher under Cournot Model so, Firm 2 will prefer Cournot Model.